# Avoiding Catastrophe in Online Learning by Asking for Help

### v4

@@ A.1 Proof roadmap
    Prove that $f(T)=\sqrt{(|Q_{T}|+1)T}$ is asymptotically between $|Q_{T}|$ and $T$ (Lemma A.1).
@@
  3. 3.
    Show that with high probability, $\sum_{j\in A}|M_{j}|$ is adequately large (Lemma A.3).
@@
    The key lemma is Lemma A.4, which shows that a randomly sampled $\boldsymbol{a}$ produces poor agent performance with high probability. The central idea is that at least $f(T)-|Q_{T}|$ sections are never queried (which is large, by Lemma A.1), so the agent has no way of knowing the optimal action in those sections. As a result, the agent picks the wrong answer at least half the time on average (and at least a quarter of the time with high probability). Lemma A.3 implies that a constant fraction of those time steps will have quite suboptimal payoffs, again with high probability.
@@
    Finally, $\sup\limits_{\boldsymbol{\mu}}\ \operatorname*{\mathbb{E}}\limits_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})\geq\operatorname*{\mathbb{E}}\limits_{\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \operatorname*{\mathbb{E}}\limits_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ R_{T}(\boldsymbol{x},\boldsymbol{y},\mu_{f,\boldsymbol{a}},\pi^{m})$, where $U(\{0,1\}^{f(T)})$ is the uniform distribution over bit strings of length $f(T)$. This is essentially an application of the probabilistic method: if a randomly chosen $\mu_{f,\boldsymbol{a}}$ has high expected regret, then the worst case $\boldsymbol{\mu}$ also has high expected regret.
While reading the proof, one should keep in mind the following:
  1. 1.
  2. 2.
    We will assume that the worst-case (not expected) number of queries is sublinear in $T$. Proposition E.2 shows that relaxing this to expected sublinear queries would not affect our conclusion.
  3. 3.
    While the intuition provided in Section 4.1 is accurate, the analysis will mostly occur in log space, so the bounds will look different. However, bounds of the form discussed in Section 4.1 can still be found as an intermediate step in Part 4 of the proof of Lemma A.4.
A.2 Proof
@@
Let $a,b:\mathbb{N}\to\mathbb{N}$ be functions such that $a(x)\in o(b(x))$. Then $c(x)=\sqrt{a(x)b(x)}$ satisfies $a(x)\in o(c(x))$ and $c(x)\in o(b(x))$.
@@
let $A\subseteq[f(T)]$ be any nonempty subset of sections. Then
$\Pr\left[\sum_{j\in A}|M_{j}|\leq\frac{T|A|}{4f(T)}\right]\leq\exp\left(\frac{-T}{16f(T)}\right)$
@@
[…]
$\Pr\left[\sum_{j\in A}|M_{j}|\leq\frac{T|A|}{4f(T)}\right]\leq\exp\left(\frac{-T|A|}{16f(T)}\right)\leq\exp\left(\frac{-T}{16f(T)}\right)$
with the last step due to $|A|\geq 1$. ∎
@@
Let $f,g:\mathbb{N}\to\mathbb{N}$. Suppose $|Q_{T}|\leq g(T)$ always and independently sample $\boldsymbol{a}\sim U(\{0,1\}^{f(T)})$ and $\boldsymbol{x}\sim U^{T}$. [^14] That is, the entire set $\{a_{1},\dots,a_{f(T)},x_{1},\dots,x_{T}\}$ is mutually independent. Then with probability at least $1-\exp\big{(}\frac{-T}{16f(T)}\big{)}-\exp\big{(}-\frac{f(T)-g(T)}{16}\big{)}$,
$\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\leq\exp\left(-\frac{LT(f(T)-g(T))}{2^{7}f(T)^{2}}\right)$
@@
Part 1: setup. Let $J_{\neg Q}=\{j\in[f(T)]:x_{t}\not\in X_{j}\ \forall t\in Q_{T}\}$ be the set of sections that are never queried. Since each query appears in exactly one section (because each input appears in exactly one section), $|J_{\neg Q}|\geq f(T)-|Q_{T}|\geq f(T)-g(T)$.
For each $j\in J_{\neg Q}$, let $y_{j}$ be the action taken most frequently among time steps in $M_{j}$:
$y_{j}=\operatorname*{arg\,max}_{y\in\{0,1\}}\Big{|}\{t\in M_{j}:y=y_{t}\}\Big{|}$
Let $\bar{J}=\{j\in J_{\neg Q}:a_{j}\neq y_{j}\}$. For each $j\in\bar{J}$, let $M_{j}^{\prime}=\{t\in M_{j}:y_{t}\neq a_{j}\}$ be the set of time steps where the agent chooses the wrong action (assuming payoff function $\mu_{f,\boldsymbol{a}}$).
[…]
$\Pr\left[|\bar{J}|\leq\frac{f(T)-g(T)}{4}\right]\leq\exp\left(-\frac{f(T)-g(T)}{16}\right)$
Part 3: $|M_{j}^{\prime}|\geq|M_{j}|/2$. Since $j\in J_{\neg Q}$, the mentor is not queried on any time step $t\in M_{j}$, so $y_{t}\in\{0,1\}$ for all $t\in M_{j}$. Since the agent chooses one of two actions for each $t\in M_{j}$, the more frequent action must be chosen chosen at least half of the time: $y_{t}=y_{j}$ for at least half of the time steps in $M_{j}$. Since $a_{j}\neq y_{j}$ for $j\in\bar{J}$, we have $y_{t}=y_{j}\neq a_{j}$ for those time steps, so $|M_{j}^{\prime}|\geq|M_{j}|/2$.
Part 4: a bound in terms of $\bar{J}$ and $M_{j}$. Consider any $j\in\bar{J}$ and $t\in M_{j}^{\prime}\subseteq M_{j}$. By definition of $M_{j}$, we have $|m_{j}-x_{t}|\leq\frac{1}{4f(T)}$. Then by definition of $\mu_{f,\boldsymbol{a}}$,
@@
$=$ $\ =1-\frac{L}{4f(T)}$
Now aggregating across time steps,
@@
$\log\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\leq$ $\ \log\prod_{j\in\bar{J}}\left(1-\frac{L}{4f(T)}\right)^{|M_{j}|/2}$
$=$ $\ \sum_{j\in\bar{J}}\frac{|M_{j}|}{2}\log\left(1-\frac{L}{4f(T)}\right)$
$\leq$ $\ -\sum_{j\in\bar{J}}\frac{L|M_{j}|}{8f(T)}$
[…]
$\log\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\leq$ $\ -\sum_{j\in\bar{J}}\frac{L|M_{j}|}{8f(T)}$
$\leq$ $\ -\frac{L}{8f(T)}\cdot\frac{T|\bar{J}|}{4f(T)}$
$\leq$ $\ -\frac{L}{8f(T)}\cdot\frac{T}{4f(T)}\cdot\frac{f(T)-g(T)}{4}$
$=$ $\ -\frac{LT(f(T)-g(T))}{2^{7}f(T)^{2}}$
Exponentiating both sides proves the lemma. ∎
Let $\alpha(T)=\exp\big{(}\frac{-T}{16f(T)}\big{)}+\exp\big{(}-\frac{f(T)-g(T)}{16}\big{)}$ for brevity (with $f$ and $g$ to be chosen later). See 4.1
Proof.
If the algorithm has sublinear queries, then there exists $g(T)\in o(T)$ such that $|Q_{T}|\leq g(T)$ always. Then by Lemma A.1, $g(T)\in o(\sqrt{g(T)T})$ and $\sqrt{g(T)T}\in o(T)$. Let $f(T)=\lceil\sqrt{g(T)T}\rceil$: then $f(T)\in\Theta(\sqrt{g(T)T})$, so $g(T)\in o(f(T))$ and $f(T)\in o(T)$. First, this implies that $\lim_{T\to\infty}\alpha(T)=0$. Next, $g(T)\in o(f(T))$ implies that exists $T_{0}$ such that $g(T)\leq f(T)/2$ for all $T\geq T_{0}$. Combining this with Lemma A.4 and noting that $\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\leq 1$, for all $T\geq T_{0}$ we have
$\ \operatorname*{\mathbb{E}}_{\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ \>\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})$
$\leq$ $\ \alpha(T)\cdot 1+\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LT(f(T)-|g(T))}{2^{7}f(T)^{2}}\right)$
$\leq$ $\ \alpha(T)+\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LTf(T)/2}{2^{7}f(T)^{2}}\right)$
$=$ $\ \alpha(T)+\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LT}{2^{8}f(T)}\right)$
Since $\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}^{m}(x_{t})=1$ always, we have [^15] Fubini’s theorem means we need not worry about the order of the expectation operators.
$\sup_{\boldsymbol{\mu}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})\geq$ $\ \operatorname*{\mathbb{E}}_{\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ R_{T}(\boldsymbol{x},\boldsymbol{y},(\mu_{f,\boldsymbol{a}},\dots,\mu_{f,\boldsymbol{a}}),\pi^{m})$
$=\operatorname*{\mathbb{E}}_{\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}$ $\ \,\left[\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}^{m}(x_{t})-\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\right]$
$\geq$ $\ 1-\alpha(T)-\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LT}{2^{8}f(T)}\right)$
Therefore
$\lim_{T\to\infty}\sup_{\boldsymbol{\mu}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})\geq$ $\ 1-\lim_{T\to\infty}\alpha(T)-\big{(}1-\lim_{T\to\infty}\alpha(T)\big{)}\cdot\exp\left(\lim_{T\to\infty}-\frac{LT}{2^{8}f(T)}\right)$
$=$ $\ 1-0-(1-0)\cdot\exp(-\infty)$
[…]
as required. ∎

### v5

@@ A.2 Proof roadmap
    Show that multiplicative regret and additive regret are tightly related (Lemma A.1). We will also use this lemma for our positive results.
@@
    Prove an asymptotic density lemma which we will use to show that $f(T)=\sqrt{|Q_{T}|T}$ is asymptotically between $|Q_{T}|$ and $T$ (Lemma A.2).
  3. 3.
@@
    Show that with high probability, $\sum_{j\in S}|M_{j}|$ is large for any subset of sections $S$ (Lemma A.4).
@@
    Prove that $|J_{\neg Q}^{\prime}|$ is large with high probability (Lemma A.5).
  6. 6.
    The key lemma is Lemma A.6, which shows that a randomly sampled $\boldsymbol{a}$ produces poor agent performance with high probability. The central idea is that at least $f(T)-|Q_{T}|$ sections are never queried (which is large, by Lemma A.2), so the agent has no way of knowing the optimal action in those sections. As a result, the agent picks the wrong answer at least half the time on average (and at least a quarter of the time with high probability). Lemma A.4 implies that a constant fraction of those time steps will have significantly suboptimal payoffs, again with high probability.
  7. 7.
    Apply $\sup\limits_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}\limits_{\boldsymbol{x},\boldsymbol{y}}\ R_{T}^{\times}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})\geq\operatorname*{\mathbb{E}}\limits_{\pi^{m},\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \operatorname*{\mathbb{E}}\limits_{\boldsymbol{x},\boldsymbol{y}}\ R_{T}^{\times}(\boldsymbol{x},\boldsymbol{y},\mu_{f,\boldsymbol{a}},\pi^{m})$. Here $U(\{0,1\}^{f(T)})$ is the uniform distribution over bit strings of length $f(T)$ and we write $\pi^{m},\boldsymbol{a}\sim U(\{0,1\}^{f(T)})$ with slight abuse of notation, since $\pi^{m}$ is not drawn from $U(\{0,1\}^{f(T)})$ but rather is determined by $\boldsymbol{a}$ which is drawn from $U(\{0,1\}^{f(T)})$.
  8. 8.
    The analysis above results in a lower bound on $R_{T}^{+}$. The last step is to use Lemma A.1 to obtain a lower bound on $R_{T}^{\times}$.
Step 7 is essentially an application of the probabilistic method: if a randomly chosen $\mu_{f,\boldsymbol{a}}$ has high expected regret, then the worst-case $\boldsymbol{\mu}$ also has high expected regret. We have included subscripts in the expectations above to distinguish between the randomness over $\boldsymbol{a}$ and $\boldsymbol{x},\boldsymbol{y}$. When subscripts are omitted, the expected value is over all randomness, i.e., $\boldsymbol{a},\boldsymbol{x},$ and $\boldsymbol{y}$.
A.3 Proof
@@
[…]
Proof.
Recall the standard inequalities $11-\frac{1}{a}\leq\log a\leq a-$ 1 for any $a>0$.
Part 1: $R_{T}^{+}\leq R_{T}^{\times}$. If $\mu_{t}(x_{t},y_{t})=0$ for any $t\in[T]$, then $R_{T}^{\times}=\infty$ and the claim is trivially satisfied. Thus assume $\mu_{t}(x_{t},y_{t})>0$ for all $t\in[T]$.
$R_{T}^{+}=$ $\ \sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})$ (Definition of $(\text{Definition of $R_{T}^{+}$})$
[…]
$=$ $\ R_{T}^{\times}$ (Definition of $(\text{Definition of $R_{T}^{\times}$})$
Part 2: $R_{T}^{\times}\leq R_{T}^{+}/\min_{t\in[T]}\mu_{t}(x_{t},y_{t})$. We have
$R_{T}^{\times}=$ $\ \log\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\log\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$ (Definition of RT× given $(\text{Definition of $R_{T}^{\times}$ given $\mu_{t}(x_{t},y_{t})>0\ \forall t\in[T]$})$
[…]
$=$ $\ \frac{R_{T}^{+}}{\min_{t\in[T]}\mu_{t}(x_{t},y_{t})}$ (Definition of $(\text{Definition of $R_{T}^{+}$})$
as claimed. ∎
Lemma A.2.
Let $a,b:\mathbb{R}_{>0}\to\mathbb{R}_{>0}$ be functions such that $a(x)\in o(b(x))$. Then $c(x)=\sqrt{a(x)b(x)}$ satisfies $a(x)\in o(c(x))$ and $c(x)\in o(b(x))$.
@@
Let $S\subseteq[f(T)]$ be any nonempty subset of sections. Then
$\Pr\left[\sum_{j\in S}|M_{j}|\leq\frac{T|S|}{4f(T)}\right]\leq\exp\left(\frac{-T}{16f(T)}\right)$
@@
Fix any $j\in[f(T)]$. For each $t\in[T]$ , define the random variable $z_{t}$ by $z_{t}=1$ if $t\in M_{j}$ for some $j\in S$ and 0 otherwise. We have $t\in M_{j}$ iff $x_{t}$ falls within a particular interval of length $\frac{1}{2f(T)}$. Since these intervals are disjoint for different $j$’s, we have $z_{t}=1$ iff $x_{t}$ falls within a portion of the input space with total measure $\frac{|S|}{2f(T)}$. Since $x_{t}$ is uniformly random across $[0,1]$, we have $\operatorname*{\mathbb{E}}[z_{t}]=\frac{|S|}{2f(T)}$. Then $\operatorname*{\mathbb{E}}[\sum_{t=1}^{T}z_{t}]=\operatorname*{\mathbb{E}}[\sum_{j\in S}|M_{j}|]=\frac{T|S|}{2f(T)}$. Furthermore, since $x_{1},\dots,x_{T}$ are i.i.d., so are $z_{1},\dots,z_{T}$. Then by Lemma A.3,
$\Pr\left[\sum_{j\in S}|M_{j}|\leq\frac{T|S|}{4f(T)}\right]\leq\exp\left(\frac{-T|S|}{16f(T)}\right)\leq\exp\left(\frac{-T}{16f(T)}\right)$
with the last step due to $|S|\geq 1$. ∎
@@
We have
$\Pr\left[|J_{\neg Q}^{\prime}|\leq\frac{f(T)-\operatorname*{\mathbb{E}}[|Q_{T}|]}{4}\right]\leq\exp\left(-\frac{f(T)-\operatorname*{\mathbb{E}}[|Q_{T}|]}{16}\right)$
@@
Define a random variable $z_{j}=\mathbf{1}_{j\in J_{\neg Q}^{\prime}}$ for each $j\in J_{\neg Q}$. By definition, if $j\in J_{\neg Q}$, no input in $X_{j}$ is queried. Since queries outside of $X_{j}$ provide no information about $a_{j}$, the agent’s actions must be independent of $a_{j}$. In particular, the random variables $a_{j}$ and $y^{j}$ are independent. Combining that independence with $\Pr[a_{j}=0]=\Pr[a_{j}=1]=0.5$ yields $\Pr[z_{j}=1]=0.5$ for all $j\in J_{\neg Q}$. Then
$\operatorname*{\mathbb{E}}\left[|J_{\neg Q}^{\prime}|\right]=$ $\ \operatorname*{\mathbb{E}}\left[\sum_{j\in J_{\neg Q}}z_{j}\right]$
$=$ $\ |J_{\neg Q}|/2$
$\geq$ $\ \frac{f(T)-\operatorname*{\mathbb{E}}[|Q_{T}|]}{2}$
Furthermore, since $a_{1},\dots,a_{f(T)}$ are independent, the random variables $\{z_{j}:j\in J_{\neg Q}\}$ are also independent. Applying Lemma A.3 yields the desired bound. ∎
Lemma A.6.
Suppose $f:\mathbb{N}\to\mathbb{N}$ and independently sample $\boldsymbol{a}\sim U(\{0,1\}^{f(T)})$ and $\boldsymbol{x}\sim U(\mathcal{X})^{T}$. [^11] That is, the entire set $\{a_{1},\dots,a_{f(T)},x_{1},\dots,x_{T}\}$ is mutually independent. Then with probability at least $1-\exp\big{(}\frac{-T}{16f(T)}\big{)}-\exp\big{(}-\frac{f(T)-\operatorname*{\mathbb{E}}[|Q_{T}|]}{16}\big{)}$,
$R_{T}^{+}\geq\frac{LT(f(T)-\operatorname*{\mathbb{E}}[|Q_{T}|])}{2^{7}f(T)^{2}}$
Proof.
Consider any $j\in J_{\neg Q}^{\prime}$ and $t\in M_{j}^{\prime}\subseteq M_{j}$. By definition of $M_{j}$, we have $|m_{j}-x_{t}|\leq\frac{1}{4f(T)}$. Then by the definition of $\mu_{f,\boldsymbol{a}}$,
@@
$=$ $\ 1-\frac{L}{4f(T)}$
Since $\mu_{f,\boldsymbol{a}}^{m}(x_{t})\geq\mu_{f,\boldsymbol{a}}(x_{t},y_{t})$ always, we can safely restrict ourselves to time steps $t\in M_{j}^{\prime}$ for some $j\in J_{\neg Q}^{\prime}$ and still obtain a lower bound:
$R_{T}^{+}=$ $\ \sum_{t=1}^{T}(\mu_{f,\boldsymbol{a}}^{m}(x_{t})-\mu_{f,\boldsymbol{a}}(x_{t},y_{t}))$ (Definition of $(\text{Definition of $R_{T}^{+}$})$
$\geq$ $\ \sum_{j\in J_{\neg Q}^{\prime}}\sum_{t\in M_{j}^{\prime}}\big{(}\mu_{f,\boldsymbol{a}}^{m}(x_{t})-\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\big{)}$ $(\mu_{f,\boldsymbol{a}}^{m}(x_{t})\geq\mu_{f,\boldsymbol{a}}(x_{t},y_{t}))$
$\geq$ $\ \sum_{j\in J_{\neg Q}^{\prime}}\sum_{t\in M_{j}^{\prime}}(1-\mu_{f,\boldsymbol{a}}(x_{t},y_{t}))$ $(\mu_{f,\boldsymbol{a}}^{m}(x_{t})=1\text{ always})$
[…]
Since $j\in J_{\neg Q}$, the mentor is not queried on any time step $t\in M_{j}$, so $y_{t}\in\{0,1\}$ for all $t\in M_{j}$. Since the agent chooses one of two actions for each $t\in M_{j}$, the more frequent action must be chosen at least half of the time: $y_{t}=y^{j}$ for at least half of the time steps in $M_{j}$. Since $a_{j}\neq y^{j}$ for $j\in J_{\neg Q}^{\prime}$, we have $y_{t}=y^{j}\neq a_{j}$ for those time steps, so $|M_{j}^{\prime}|\geq|M_{j}|/2$. Thus
$R_{T}^{+}\geq\sum_{j\in J_{\neg Q}^{\prime}}\frac{L|M_{j}|}{8f(T)}$
[…]
$R_{T}^{+}\geq$ $\ \sum_{j\in J_{\neg Q}^{\prime}}\frac{L|M_{j}|}{8f(T)}$
$\geq$ $\ \frac{L}{8f(T)}\cdot\frac{T|J_{\neg Q}^{\prime}|}{4f(T)}$
$\geq$ $\ \frac{L}{8f(T)}\cdot\frac{T}{4f(T)}\cdot\frac{f(T)-\operatorname*{\mathbb{E}}[|Q_{T}|]}{4}$
$=$ $\ \frac{LT(f(T)-\operatorname*{\mathbb{E}}[|Q_{T}|])}{2^{7}f(T)^{2}}$
as required. ∎
For a given $f:\mathbb{N}\to\mathbb{N}$, define $\alpha_{f}(T)=\exp\big{(}\frac{-T}{16f(T)}\big{)}+\exp\big{(}-\frac{f(T)-\operatorname*{\mathbb{E}}[|Q_{T}|]}{16}\big{)}$ for brevity. See 4.1
Proof.
If the algorithm has sublinear queries, then there exists $g(T)\in o(T)$ such that $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}[|Q_{T}|]\leq g(T)$. Consider any such $g(T)$ satisfying $g(T)>0$. Since this holds for every $\boldsymbol{\mu}$, it also holds in expectation over $\boldsymbol{a}\sim U(\{0,1\})^{f(T)}$, so $\operatorname*{\mathbb{E}}_{\boldsymbol{a},\boldsymbol{x},\boldsymbol{y}}[|Q_{T}|]=\operatorname*{\mathbb{E}}[|Q_{T}|]\leq g(T)$.
Next, Lemma A.2 gives us $g(T)\in o(\sqrt{g(T)T})$ and $\sqrt{g(T)T}\in o(T)$. Let $f(T)=\lceil\sqrt{g(T)T}\rceil$: then $f(T)\in\Theta(\sqrt{g(T)T})$, so $g(T)\in o(f(T))$ and $f(T)\in o(T)$. First, this implies that $\lim_{T\to\infty}\alpha_{f}(T)=0$. Second, $g(T)\in o(f(T))$ implies that exists $T_{0}$ such that $g(T)\leq f(T)/2$ for all $T\geq T_{0}$. We also have $R_{T}^{+}\geq 0$ always since $\mu_{f,\boldsymbol{a}}^{m}(x_{t})\geq\mu_{f,\boldsymbol{a}}(x_{t},y_{t})$ always. Then for all $T\geq T_{0}$ we have
$\ \operatorname*{\mathbb{E}}_{\pi^{m},\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[R_{T}^{+}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})\right]$
@@
Since $\lim_{T\to\infty}\alpha_{f}(T)=0$,
$\sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \big{[}R_{T}^{+}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})\big{]}\geq$ $\ \operatorname*{\mathbb{E}}_{\pi^{m},\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \big{[}R_{T}^{+}(\boldsymbol{x},\boldsymbol{y},(\mu_{f,\boldsymbol{a}},\dots,\mu_{f,\boldsymbol{a}}),\pi^{m})\big{]}$
$\geq$ $\ \big{(}1-\alpha_{f}(T)\big{)}\left(\frac{LT}{2^{8}f(T)}\right)$
$\in$ $\ \Omega\left(\frac{LT}{f(T)}\right)$
$=$ $\ \Omega\left(L\sqrt{\frac{T}{g(T)}}\right)$
This holds for any $g(T)\in o(T)$ such that $\sup_{\boldsymbol{\mu}}\operatorname*{\mathbb{E}}[|Q_{T}|]\leq g(T)$ and $g(T)>0$. Thus we can simply set $g(T)=\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}|]+1$, since $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}|]$ is indeed a function of only $T$.
Since $\mu_{f,\boldsymbol{a}}^{m}(x_{t})\geq\mu_{f,\boldsymbol{a}}(x_{t},y_{t})$ for all $t\in[T]$, Lemma A.1 implies that
$\sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \big{[}R_{T}^{\times}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})\big{]}\geq$ $\ \operatorname*{\mathbb{E}}_{\pi^{m},\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \big{[}R_{T}^{\times}(\boldsymbol{x},\boldsymbol{y},(\mu_{f,\boldsymbol{a}},\dots,\mu_{f,\boldsymbol{a}}),\pi^{m})\big{]}$
$\in$ $\ \Omega\left(L\sqrt{\frac{T}{\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}|]+1}}\right)$
completing the proof.∎
See 4.1.1
Proof.
We have $\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}^{m}(x_{t})=1$ from our construction. Then Lemma A.1 implies that $R_{T}^{\times}\geq R_{T}^{+}$, so
$\exp(-R_{T}^{+})\geq$ $\ \exp(-R_{T}^{\times})$
$=$ $\ \exp\left(\log\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})-\log 1\right)$
$=$ $\ \prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})$
Then by Lemma A.6, with probability $1-\alpha_{f}(T)$,
$\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\leq\exp\left(-\frac{LT(f(T)-\operatorname*{\mathbb{E}}[|Q_{T}|])}{2^{7}f(T)^{2}}\right)\leq\exp\left(-\frac{LT}{2^{8}f(T)}\right)$
Since $\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\leq 1$ always,
$\lim_{T\to\infty}\operatorname*{\mathbb{E}}_{\pi^{m},\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\right]\leq$ $\ \lim_{T\to\infty}\operatorname*{\mathbb{E}}_{\pi^{m},\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\right]$
$\leq$ $\ \lim_{T\to\infty}\left(1\cdot\alpha_{f}(T)+(1-\alpha_{f}(T))\cdot\exp\left(-\frac{LT}{2^{8}f(T)}\right)\right)$
$\leq$ $\ \lim_{T\to\infty}(1-\alpha_{f}(T))\cdot\lim_{T\to\infty}\exp\left(-\frac{LT}{2^{8}f(T)}\right)$
$=$ $\ 1\cdot\exp(-\infty)$
@@
Since this upper bound holds for a randomly chosen $\mu_{f,\boldsymbol{a}},\pi^{m}$, the same upper bound holds for a worst-case choice of $\boldsymbol{\mu},\pi^{m}$ among $\boldsymbol{\mu},\pi^{m}$ which satisfy $\mu_{t}^{m}(x)=1$ for all $t\in[T],x\in\mathcal{X}$. Formally,
$\lim_{T\to\infty}\ \sup_{\boldsymbol{\mu},\pi^{m}:\,\mu_{t}(x)=1\,\forall t,x}\operatorname*{\mathbb{E}}\left[\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})\right]\leq 0$
Since $\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})\geq 0$ always, the inequality above holds with equality. ∎